OpenStudy (anonymous):

need help verifying answers find the zeros of : f(x)=x^2 (2e^2x) +(2xe^2x) + (e^2x) + (2xe^2x) I get : x= -1+ (1/2)sqrt 2 and x=-1-(1/2)sqrt 2

OpenStudy (blurbendy):

x = 0 x = -1 -1/Sqrt[2] x = 1/Sqrt[2] - 1

OpenStudy (anonymous):

i dont distribute he x^2 therefore x=0 and the other two change as well?

OpenStudy (blurbendy):

wait, what?

OpenStudy (anonymous):

\[f(x)= x^2 (2e^{2x})+2xe ^{2x}+e ^{2x}+2xe ^{2x}\]

OpenStudy (blurbendy):

Yes, i know. trust me, my answer is correct. plug the values in if you don't believe me.

OpenStudy (anonymous):

yea i trust you, my question is you set all of that equal to 0. i distributed the x^2 did you?

OpenStudy (anonymous):

take e^2x common \[\Large f(x)=2e^{2x}(x^2+2x+1/2)\] set it equal to zero \[\Large 2e^{2x}=0\] no solution can be obtaind for this . so solve \[\Large x^2+2x+1/2=0\] by completing square x^2+2x =-1/2 x^2+2x+1=1/2 (x+1)^2=1/2 take square root of both sides \[\Large x=-1-\frac{1}{\sqrt{2}}\] \[\Large x=\frac{1}{\sqrt{2}}-1\] these are the possible zeros .

OpenStudy (anonymous):

in the first step take 2e^2x common !